OpenStudy (anonymous):

L'Hopitals rule: evaluate the limit:

OpenStudy (amistre64):

lhops rule works for limits that give us 0/0 or inf/inf, if memory serves

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ \sin^2 x }{ \tan(x^2) }\]

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

so the first thing i think I'd do is take the derivative of the top and the bottom, because if i just try plugging in 0 i'll get 0 in the denominator

OpenStudy (solomonzelman):

you can apply the L'H'S now

OpenStudy (solomonzelman):

Hints: \(\large\color{black}{\sin^2x=(\sin~x)^2}\) and then use the chain rule. for \(\large\color{black}{\tan(x^2)}\) you will need a chain rule also.

OpenStudy (solomonzelman):

hey guys, let him try

OpenStudy (anonymous):

right, so i got. \[\frac{2 \sin x \cos x}{2xsec^2 x^2}\]

OpenStudy (jhannybean):

\[\sin^2(x) +\cos^2(x) = 1\]For \(\sin^2(x)\) , can you not replace it with \(1-\cos^2(x)\)?

OpenStudy (zarkon):

L'Hospitals rule is overkill for this problem

OpenStudy (solomonzelman):

the derivative is correct, study guy

OpenStudy (anonymous):

okay.

OpenStudy (solomonzelman):

I would like to hear another way to do it asides from L'h'S if there is any. I am not good at mathematics, and I might miss a better approach sometimes.

OpenStudy (nincompoop):

hospitalization may be overkill, but I think the purpose of the question is to get someone used to the concept

OpenStudy (jhannybean):

hospitalization LOL

OpenStudy (anonymous):

okay so you can replace the sin^2 x with 1 - cos^2 x because of trig identites

OpenStudy (solomonzelman):

I mean, it sounds like you HAVE TO use it, so we will go that way.

OpenStudy (zarkon):

I just wish they gave problems to students that necessitated the need for L'Hospitals rule

OpenStudy (jhannybean):

Yeah, before taking derivatives I would always reference trig identities first so you can simplify the function before using any simplification rules.

OpenStudy (nincompoop):

it's like asking someone to evaluate using limit definition... it is overkill, but it serves a purpose

OpenStudy (solomonzelman):

Well, anyway, it is a requirement so lets proceed. We got: \(\Large\color{black}{\displaystyle\lim_{x \rightarrow ~a} ~\frac{ 2\sin(x)\cos(x)}{ 2x\sec^2(2x)}}\)

OpenStudy (solomonzelman):

Okay, when you plug in I mean not a, \(\Large\color{black}{\displaystyle\lim_{x \rightarrow ~0} ~\frac{ 2\sin(x)\cos(x)}{ 2x\sec^2(2x)}}\)

OpenStudy (solomonzelman):

if you plug in zero into the limit, what do you get, study guy?

OpenStudy (anonymous):

well the denominator is 0 again. so we use lhopital's again?

OpenStudy (solomonzelman):

No no

OpenStudy (solomonzelman):

if the denominator and numerator is zero.

OpenStudy (solomonzelman):

and this is the case, but I just want to tell the concept, that if you get (for example) 1/0, you can NOT apply L'H'S.

OpenStudy (solomonzelman):

but in this case it is still 0/0, right because sin(0) is zero, and the 2x is zero, when x=0.

OpenStudy (solomonzelman):

\(\Large\color{black}{\displaystyle\lim_{x \rightarrow ~0} ~\frac{ 2\sin(x)\cos(x)}{ 2x\sec^2(2x)}}\)

OpenStudy (solomonzelman):

Apply L'H's rule again.

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

take your time:)

OpenStudy (anonymous):

should i first use the trig identity to make the 2sin(x)cos(x) into sin(2x) so it's easier?

OpenStudy (jhannybean):

Yes!!

OpenStudy (solomonzelman):

yes, if you want.

OpenStudy (solomonzelman):

\(\Large\color{black}{2\sin(x)\cos(x)=\sin(x)\cos(x)+\sin(x)\cos(x)=?}\)

OpenStudy (jhannybean):

@SolomonZelman keep it simple!!

OpenStudy (solomonzelman):

Am I not?

OpenStudy (jhannybean):

Lol.. taking the derivative of \(2x\sec^2(2x)\) is also a headache enough.

OpenStudy (anonymous):

okay so i think i got it is it: \[\frac{ 2\cos(2x )}{ 4\sec(x^2)\sec(x^2)\tan(x^2) }\]

OpenStudy (solomonzelman):

not really. taking a derivativ is never a headache. A couple of chains, even I can do that.

OpenStudy (zarkon):

\[\Large\color{black}{\displaystyle\lim_{x \rightarrow ~0} ~\frac{ 2\sin(x)\cos(x)}{ 2x\sec^2(2x)}}\] \[=\Large\color{black}{\displaystyle\lim_{x \rightarrow ~0} ~\frac{ \sin(x)}{ x}\lim_{x \rightarrow ~0}\frac{\cos(x)}{\sec^2(2x)}}\]

OpenStudy (jhannybean):

^ That's what I was getting at.

OpenStudy (anonymous):

oh yeah because of the product rule of limits.

OpenStudy (jhannybean):

And before learning LH rule, you should already know what the identity \(\dfrac{\sin(x)}{x}\) gives you.

OpenStudy (solomonzelman):

yes, but I thought LHS was required.

OpenStudy (zarkon):

you already used L'Hospitals rule once

OpenStudy (jhannybean):

Sometimes when you are able to spot an identity, you can avoid using LH rule.

OpenStudy (solomonzelman):

true:) But you know how some teacher insist on re-using it again and again whether it's easy or not... although your way really seems convincing and (almost) the only...

OpenStudy (jhannybean):

The purpose of LH rule is to simplify a function and bring it out of an indeterminant form. Once it is out of indeterminant form,you can do whatever you want. It's a free country.

OpenStudy (jhannybean):

:P

OpenStudy (solomonzelman):

sure:)

OpenStudy (anonymous):

we're left with cos x over sec squared 2x since sinx over x is 1 right?

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

\(\Large\color{black}{\displaystyle\lim_{x \rightarrow ~0} ~\frac{ 2\sin(x)\cos(x)}{ 2x\sec^2(2x)}}\) \(\Large\color{black}{\displaystyle\lim_{x \rightarrow ~0} ~\frac{ \sin(x)\cos(x)}{ x\sec^2(2x)}}\) \(\Large\color{black}{\displaystyle\lim_{x \rightarrow ~0} ~\frac{ \sin(x)}{ x}\times }\)\(\Large\color{black}{\displaystyle\lim_{x \rightarrow ~0} ~\frac{ \cos(x)}{ \sec^2(2x)}}\)

OpenStudy (solomonzelman):

this is what we have done.

OpenStudy (solomonzelman):

And then, as you said. \(\Large\color{black}{1\times }\)\(\Large\color{black}{\displaystyle\lim_{x \rightarrow ~0} ~\frac{ \cos(x)}{ \sec^2(2x)}}\)

OpenStudy (anonymous):

so now i j just plug in 0 right?

OpenStudy (solomonzelman):

yes, everything is correct. I am just again reposting the work, just to make sure you have all the information and stpes you need to have.

OpenStudy (solomonzelman):

yes, plug in 0

OpenStudy (anonymous):

so i got 1 as my limit.

OpenStudy (solomonzelman):

Yup!

OpenStudy (solomonzelman):

the answer is 1

OpenStudy (jhannybean):

Good job.

OpenStudy (zarkon):

I still think it is better to do the following \[\frac{ \sin^2 x }{ \tan(x^2) }\] \[=\frac{\sin(x)}{x}\frac{\sin(x)}{x}\frac{x^2}{\tan(x^2)}\] \[=\frac{\sin(x)}{x}\frac{\sin(x)}{x}\cos(x^2)\frac{x^2}{\sin(x^2)}\to1\times 1\times1\times1=1\]

OpenStudy (solomonzelman):

it's definitely IS better to do that.

OpenStudy (solomonzelman):

anyways, just a little attachment.

OpenStudy (zarkon):

save L'Hospitals rule for hard problems

OpenStudy (jhannybean):

Ugh, accidentally deleted my post.

OpenStudy (zarkon):

\[\tan(x^2)=\frac{\sin(x^2)}{\cos(x^2)}\]

OpenStudy (zarkon):

\[\lim_{x\to 0}\frac{\sin(x^2)}{x^2}=\lim_{u\to 0}\frac{\sin(u)}{u}=1\] \[u=x^2\]

OpenStudy (jhannybean):

oh I got it.

OpenStudy (solomonzelman):

I think it comes not from substitution that \(\large\color{black}{\displaystyle\lim_{x \rightarrow ~0}\frac{\sin^2 x}{x^2}=0}\). \(\large\color{black}{\displaystyle\lim_{x \rightarrow ~0}\left(\begin{matrix} \frac{\sin x}{x} \\ \end{matrix}\right)^2=0}\) \(\large\color{black}{\left(\begin{matrix}\displaystyle\lim_{x \rightarrow ~0} \frac{\sin x}{x} \\ \end{matrix}\right)^2=0}\)

OpenStudy (solomonzelman):

just by applying basic limit rules. I meean sub is shorter, but rules are more basic.

OpenStudy (solomonzelman):

either way...

OpenStudy (zarkon):

those limits are 1 not 0

OpenStudy (solomonzelman):

yes 1... by bad, I left a zero from the latex that I copied, sorry.

OpenStudy (solomonzelman):

(but you get what I was trying to say)

OpenStudy (anonymous):

yeah,this makes more sense thank you all so much for your help.

OpenStudy (solomonzelman):

yw

OpenStudy (solomonzelman):

ohh, my pdf has a typo in it. where the integral is. I'll fix that, hold.

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

there fixed it.

OpenStudy (solomonzelman):

it had a and b instead of z and w, not that it matters, but .... it was a nice post, and tnx everyone!

OpenStudy (anonymous):

thanks.

OpenStudy (jhannybean):

There is an identity for cos(x) that I do not remember, it complements \[\lim_{x \rightarrow 0} \frac{\sin(x)}{x}=1\] Something like.. \[\lim_{x \rightarrow 0} \frac {1-\cos(x)}{x}=0\] I think..

OpenStudy (jhannybean):

Anyway, good luck with Calculus!!

OpenStudy (anonymous):

i'll try finding it in my book and thank you

OpenStudy (solomonzelman):

yes, correct, Jhanny